# EpochCore: Digital Hardware Accelerator For Structured State-Space Models

### v1

@@ III-B1 Traditional PE vs LIMA-PE
where $B_{s}\mapsto\mathbf{\overline{A}}$: The fixed recurrent coefficient, $A_{m}\mapsto\mathbf{\overline{B}}\cdot u(t)$: The scaled input propagating through the SA in the North-to-South direction, $Y_{s}\mapsto\mathbf{x}(t)$: The current state vector (RHS), and $Y_{s}\mapsto\mathbf{x}(t+\Delta t)$: The next state vector (LHS). This operation effectively performs recurrent integration. The MAC operation is implemented by incorporating multiplexers (M1, M2 and M5) and using cached Control Out signals to configure these multiplexers, as illustrated in Figure 6a. Traditional PEs (such as TPUs) do not support such MAC operation.
@@
where $(B_{s}+A_{m})\mapsto(\mathbf{\overline{A}}+\mathbf{\overline{B}}\cdot u(t))$: The time-varying coefficient for the recurrent integration operation. This operation is implemented by incorporating an additional full-adder and the multiplexer M7, with the relevant circuitry highlighted in blue, as shown in Figure 6b. The inclusion of M7 introduces a novel mechanism to feed the scaled input into both the multiplier coefficient and the accumulator.
The TRI-MAC operation employs a dataflow approach similar to FRI-MAC, ensuring compatibility while supporting the time-dependent dynamics of Liquid-S4 models.
@@
Novelty in the LIMA-PE design: The FRI-MAC and TRI-MAC operations allow a single-cycle computation of a recurrent element-wise vector multiplication. A row of LIMA-PE performing FRI-MAC or TRI-MAC operations compute and store the internal-state vector of the Structured-SSMs for S4 and Liquid-S4 layers. LIMA-PE ’s support for both real and complex-valued data broadens EpochCore ’s applicability to a wider range of SSM models.
Each LIMA-PE integrates a programmable clock controller within its compute and load buffer units, enabling mode-aware scheduling through clock gating of decoupled preload and compute phases, including a sleep mode that fully disables both units.
@@ III-C ProDF dataflow
A unified accelerator is essential for efficiently supporting diverse workloads. Optimal dataflows vary depending on the model type—such as S4 and Liquid-S4—and are also influenced by the dimensions of the weight, input, or output matrices in traditional GEMM-based computations. Previous approaches to efficiently support multiple dataflows have either relied on combining accelerators with distinct dataflows, as demonstrated in [36], or on modifying interconnection routes between PEs using programmable controls [30, 4]. Both strategies demand significant hardware design complexity. For example, the proposal in [30] requires the additional reorder reduction switch circuitry, introducing a 15% area overhead compared to the proposed dataflow. Furthermore, this extra circuitry is not logically adjacent to the processing elements (PEs), causing additional delays due to complex routing.
@@
Layer I - Efficient Scalar-Vector Multiplication for S4s: In S4 and Liquid-S4 models, we scale the current input by a vector of coefficients (as illustrated in Layer I of Figure 2). To efficiently perform this task within a 2D SA, a single row of LIMA-PEs can be used in parallel, enabling a 1-cycle scalar-vector multiplication. This requires the LIMA-PEs to be configured for a Banded-Matrix WS MAC operation, as depicted in Figure 6c. Unlike traditional WS MAC, where inputs and results flow in horizontal and vertical directions, respectively, our method feeds inputs diagonally into the SA and propagates partial results vertically downward to the next step of the Structured-SSM model.
@@
Unified SA computation for Layer-I and Layer-II of S4s: The compact arrangement of processing elements (PEs) optimized for both Layer I and Layer II computations enables a unified approach within a single 2D SA structure. In a traditional SA, scalar-vector products, recurrent integration, and matrix multiplication are executed as separate operations, incurring additional overhead from intermediate read/write operations to on-chip SRAM when transferring outputs between steps.
The proposed dataflow eliminates this overhead by allowing all three operations to be seamlessly performed across consecutive rows of the 2D SA. This integration significantly enhances performance, resource utilization, and energy efficiency, streamlining the computation of S4 and Liquid-S4 models.
Following are the key innovations with ProDF compared to Eyeriss [4] and FusedCNN [2]. The row-specialized pipeline in EpochCore is explicitly optimized for temporal and recurrent structure, enabling efficient mapping of SSMs, unlike the homogeneous tiles in Eyeriss and FusedCNN. EpochCore ’s programmable MACs are tailored for the algebraic diversity of structured SSMs, while conventional accelerators hardwire general-purpose MACs. Fully in-situ SSM execution in EpochCore allows recurrent models to be evaluated with zero intermediate storage overhead, which neither Eyeriss nor FusedCNN can achieve due to their layer-based compute flow.
@@ III-D Mapping S4 to EpochCore
In this section, we present how we leverage the innovations in LIMA-PE microarchitecture and ProDF to efficiently map S4 and Liquid-S4 to EpochCore.
@@ III-D1 Setting PE Modes
As outlined in the section II-A, S4 models process each element of the input sequence to generate $H$ continuous output sequences through Scalar-Vector Multiplication, Recurrent Integration, and Matrix Multiplication Figure 7b illustrates the various PE operating modes in the 2-D SA for evaluating an S4 network with a state-map size of $N=3$ and a single output sequence ($H=1$). Each element of the input sequence is fed to the first row of PEs at every clock cycle. The PEs in the first row are configured to perform scalar-vector multiplication using BWS MAC modes. The input sequence is applied diagonally to the first row of PEs, which propagates scaled values downward to the next row. The second row executes recurrent integration, updating the state-map vector corresponding to the input sequence. The remaining PE rows perform matrix multiplication, configured to use Pass Through, BWS MAC or Sleep modes as needed.
@@ III-D2 Dataflow
Next we describe the data flow for S4 network as shown in Figure 7b. For S4 layers with state-map size of $N$, the SA should be of size $(N+2)\times(N+1)$. During the Pre-Load Phase, a vector of $N+2$ values, consisting of weights and PE control signals, is read from the on-chip weight SRAM and loaded into the first column of $N+2$ PEs. Over the subsequent $N+1$ cycles, all the $(N+2)\times(N+1)$ PE tiles are preloaded with these weights and control values. In the Compute Phase, (Figure 7b), the input sequence is fed from the on-chip input/output SRAM to the first row of PEs. The first data element of the output sequence becomes available after $N+2$ cycles, and subsequent outputs are generated every clock cycle. The evaluation order is illustrated in Figure 8a. Once processing of a single input sequence is complete, the SA resets the cached state-map and partial results of all PEs, while retaining the cached weights. However, both weights and partial results must be reset between input batches.
@@ III-E Mapping Liquid-S4 to EpochCore
As explained in Section II-A1, the Liquid-S4 network shares computational similarities with the S4 network, and all EpochCore innovations for the S4 network are also relevant to the Liquid-S4. One key difference is in the recurrent integration step. Liquid-S4 introduces a time-varying, input-dependent coefficient, which requires the recurrent integration PEs to operate in TRI-MAC mode. The dataflow of the Liquid-S4 network largely mirrors that of the S4 network (as illustrated in Figure 7b). The main difference is that the second-row PEs are preconfigured to execute the TRI-MAC operation. Figure 8b highlights the specific operations of the Liquid-S4 network.
@@ III-F1 CNN/RNN/Transformers
EpochCore can be configured to execute a variety of neural network operations, such as GEMM for CNN, RNN and Transformers, by supporting common dataflows like WS, IS and OS. Figure 9 demonstrates how EpochCore can be utilized for a DNN operation using the OS dataflow. In this configuration, all LIMA-PE are preset to perform the Traditional OS (TOS) MAC operation (Figure 6d) by proper selection of multiplexers (M2, M3, M4 and M6). The inputs and weight vectors are read every cycle and fed to the SA. The input data flows in the West-to-East direction, and the weights flow in the North-to-South direction. At the end of the GEMM operation, the entire output matrix is read out. The IS and WS dataflows can be similarly implemented. Transformers predominantly use GEMM operations as well, and their applications are complementary to those of SSMs. So, transformers can be executed on EpochCore. The overhead of evaluating GEMM operations on EpochCore is shown in the second paragraph under Section IV-C1.
@@ III-G EpochCore for other SSMs
Patro et. al [25] provide a taxonomy of recently proposed SSMs, categorizing them based on their structural, gated and recurrent characteristics. Within the structured category, models such as H3 [6] implement two layers of structured SSMs: one performing a shift operation, and another utilizing a diagonal state matrix. The shift-SSM layer in such models can be implemented within EpochCore’s SA as a single row of LIMA-PE by including additional sleep PE at the beginning of the row and shifting the row to the right. SSMs in the gated category improve contextualization, as seen in GSS [23]. Mapping GSS models onto EpochCore involves evaluating the associated linear layers by appropriately setting the mode of LIMA-PE, followed by the evaluation of an on-chip non-linear activation function to implement gating. The Mamba model [7, 21], shares key characteristics with Liquid-S4 [13, 12], including the use of input-dependent time-varying state-matrix. Mamba also builds upon the framework introduced in H3 [6] and GSS [23]. In Mamba, the discretized coefficients $\mathbf{\overline{A}}$ and $\mathbf{\overline{B}}$ vary with input, while the discretization step $\triangle t$ effectively serves as a gating mechanism. Supporting Mamba models on EpochCore requires pre-loading the inputs to compute input-dependent discretized coefficients on the host. These updated coefficients must be loaded onto EpochCore as weight matrices, introducing additional overhead for input processing and data transfer. This overhead is roughly $N\times$ the performance cost of S4, where $N$ is the state-map size, making EpochCore highly inefficient. To efficiently support Mamba, two modifications to EpochCore are required. (1) Support variable discretization ($\Delta t$) via additional LIMA-PE rows, (2) new MAC operations in LIMA-PE to support gated element-wise multiplication. These modifications to EpochCore, to make it generic, are part of our future work.
@@ III-H Other Components
A complete neural network model involves additional mathematical operations at the output of each layer. EpochCore integrates on-chip specialized circuitry to support a range of non-linear activation functions and layer normalization, as shown in the EpochCore micro-architecture Figure 4b. The design details of these components will be presented in future work.
@@ IV Evaluation
In this section, we compare EpochCore accelerator against other SoTA accelerators. The evaluation results apply to both inference and training. Training of neural models requires a forward pass and a backward pass, while inference involves a forward pass. The backward pass includes two GEMM operations for weight updates. In this paper, we evaluate EpochCore for inference i.e. the forward pass using 32-bit fixed-point precision. Improving S4’s forward pass performance benefits training as well.
@@ IV-A Methodology
We designed LIMA-PE at the RTL level and synthesized it using Free-PDK45 [31], at 45nm technology node, along with NanGate’s standard cell library. For synthesis, we utilized Cadence’s Genus™ tool [1] to meet a target clock frequency of 700MHz. We created test benches and simulated 1 million cycles using Cadence’s NC-Sim™, generating a Value Change Dump (VCD) file. This file was then fed back into Genus™ to evaluate the quality of results (QoR) metrics such as power, area, and performance. To assess the figure of merit (FOM) for the LIMA-PE, we compared it against other PEs from the literature. We accounted for differences in technology nodes by applying appropriate technology scaling techniques [28].
@@
To evaluate the energy efficiency and performance of EpochCore, we developed a custom cycle simulator based on ScaleSim [26] to estimate key metrics such as cycle count, throughput, and energy consumption for various configurations. For evaluating S4-based models, we used model parameters cited in [9]. Our analysis is based on LRA datasets (Table I), where S4-based neural models remain SoTA. Additionally, we created a memory bandwidth simulator to measure continuous bandwidth requirements of EpochCore and other accelerators under LRA workloads. The memory access time, using the same technology i.e. 45 nm as the PE design, was determined using a CACTI-based memory compiler [20].
TABLE III: Evaluation of PEs. Traditional-PE, STPU-PE and four types of LIMA-PE for FixedPoint32 and Int8 operation
@@ IV-B LIMA-PE vs other PEs
Sparse-TPU (STPU) [14] was designed to handle GEMM with large sparsity, featuring a PE that supports special modes. When comparing LIMA-PE to STPU-PE, LIMA-PE demonstrates a smaller area footprint for both the 8-bit and 32-bit versions. However, for power consumption, LIMA-PE is more energy-efficient than STPU-PE in the 8-bit version but consumes more power in the 32-bit version. Sparse-TPU design in [14] used 28nm technology. Appropriate scaling was used [28] to map the power and area numbers to the 45nm node used to design LIMA-PE.
@@
[…]
@@ IV-C EpochCore vs Systolic Arrays
S4 and Liquid-S4 models involve specialized operations, such as scalar-vector product and recurrent integrations that have sparsity and so they are not well suited to be run on regular SA. To evaluate EpochCore for these models, we compare its performance against other SA based architectures such as the Sparse-TPU [14] and Boriakoff’s [3] FFT based SA accelerator. The datasets used in this evaluation were selected based on software evaluations of S4 and Liquid-S4 done by Gu et al [9], which spans vision, audio, and other challenging LRA tasks commonly used to benchmark state-of-the-art model accuracy [21]. For our experiments, we focused on a single Liquid-S4 layer with a state-map size of $N=64$, and varied sequence lengths, $T$ as shown in Table I.
Figure 10: Energy/Inference and Latency/Inference for various LRA datasets shown in Table I for Liquid-S4 model when using Boriakoff based 1-D SA [3], Sparse 2D SA and EpochCore.

### v2

@@ III-B1 Traditional PE vs LIMA-PE
where $B_{s}\mapsto\mathbf{\overline{A}}$: the fixed recurrent coefficient, $A_{m}\mapsto\mathbf{\overline{B}}\cdot u(t)$: the scaled input propagating through the SA in the North-to-South direction, $Y_{s}\mapsto\mathbf{x}(t)$: the current state vector (RHS), and $Y_{s}\mapsto\mathbf{x}(t+\Delta t)$: the next state vector (LHS). This operation effectively performs recurrent integration. The MAC operation is implemented by incorporating multiplexers (M1, M2 and M5) and using buffered Control Out signals to configure these multiplexers, as illustrated in Figure 6a. Traditional PEs (such as TPUs) do not support such MAC operation.
@@
where $(B_{s}+A_{m})\mapsto(\mathbf{\overline{A}}+\mathbf{\overline{B}}\cdot u(t))$: the time-varying coefficient for the recurrent integration operation. This operation is implemented by incorporating an additional full-adder and the multiplexer M7, with the relevant circuitry highlighted in blue, as shown in Figure 6b. The inclusion of M7 introduces a novel mechanism to feed the scaled input into both the multiplier coefficient and the accumulator. The TRI-MAC operation employs a dataflow approach similar to FRI-MAC, ensuring compatibility while supporting the time-dependent dynamics of Liquid-S4 models.
@@
Novelty in the LIMA-PE design: The FRI-MAC and TRI-MAC operations allow a single-cycle computation of a recurrent element-wise vector multiplication. A row of LIMA-PE performing FRI-MAC or TRI-MAC operations compute and store the internal-state vector of the Structured-SSMs for S4 and Liquid-S4 layers. LIMA-PE ’s support for both real and complex-valued data broadens EpochCore ’s applicability to a wider range of SSM models. Each LIMA-PE integrates a programmable clock controller within its compute and load buffer units, enabling mode-aware scheduling through clock gating of decoupled preload and compute phases, including a sleep mode that fully disables both units.
@@ III-C ProDF dataflow
A unified accelerator is essential for efficiently supporting diverse workloads. Optimal dataflows vary depending on the model type—such as S4, Liquid-S4 and DNN—and are also influenced by the dimensions of the weight, input, or output matrices. Previous approaches to efficiently support multiple dataflows have either relied on combining accelerators with distinct dataflows, as demonstrated by Xu et al [36], or on modifying interconnection routes between PEs using programmable controls as shown by Tong et al [30] and Chen et al [4]. Both strategies demand significant hardware design complexity. For example, the proposal by Tong et al  [30] requires the additional reorder reduction switch circuitry, introducing a 15% area overhead compared to the proposed dataflow. Furthermore, this extra circuitry is not logically adjacent to the processing elements (PEs), causing additional delays due to complex routing.
@@
Layer I - Efficient Scalar-Vector Multiplication: In S4 and Liquid-S4 models, we scale the current input by a vector of coefficients (as illustrated in Layer I of Figure 2). To efficiently perform this task within a 2D SA, a single row of LIMA-PEs can be used in parallel, enabling a 1-cycle scalar-vector multiplication. This requires the LIMA-PEs to be configured for a Banded-Matrix WS MAC operation, as depicted in Figure 6c. Unlike traditional WS MAC, where inputs and results flow in horizontal and vertical directions, respectively, our method feeds inputs diagonally into the SA and propagates partial results vertically downward to the next step of the Structured-SSM model.
@@
Unified SA computation for Layer-I and Layer-II of S4s: The compact arrangement of processing elements (PEs) optimized for both Layer I and Layer II computations enables a unified approach within a single 2D SA structure. In a traditional SA, scalar-vector products, recurrent integration, and matrix multiplication are executed as separate operations, incurring additional overhead from intermediate read/write operations to on-chip SRAM when transferring outputs between steps. The proposed dataflow eliminates this overhead by allowing all three operations to be seamlessly performed across consecutive rows of the 2D SA. This integration significantly enhances performance, resource utilization, and energy efficiency, streamlining the computation of S4 and Liquid-S4 models.
Following are the key innovations in ProDF compared to Eyeriss [4] and FusedCNN [2]. The row-specialized pipeline in EpochCore is explicitly optimized for temporal and recurrent structure, enabling efficient mapping of SSMs, unlike the homogeneous tiles in Eyeriss and FusedCNN. EpochCore ’s programmable MACs are tailored for the algebraic diversity of structured SSMs, while conventional accelerators hardwire general-purpose MACs. Fully in-situ SSM execution in EpochCore allows recurrent models to be evaluated with zero intermediate storage overhead, which neither Eyeriss nor FusedCNN can achieve due to their layer-based compute flow.
@@ III-D Mapping S4 to EpochCore
In this section, we present how we leverage the innovations in LIMA-PE microarchitecture and ProDF to efficiently map S4 to EpochCore.
@@ III-D1 Setting PE Modes
As outlined in the section II-A, S4 models process each element of the input sequence to generate $H$ continuous output sequences through Scalar-Vector Multiplication, Recurrent Integration, and Matrix Multiplication. Figure 7b illustrates the various PE operating modes in the 2-D SA for evaluating an S4 network with a state-map size of $N=3$ and a single output sequence ($H=1$). Each element of the input sequence is fed to the first row of PEs at every clock cycle. The PEs in the first row are configured to perform scalar-vector multiplication using BWS MAC modes. The input sequence is applied diagonally to the first row of PEs, which propagates scaled values downward to the next row. The second row executes recurrent integration, updating the state-map vector corresponding to the input sequence. The remaining PE rows perform matrix multiplication, configured to use Pass Through, BWS MAC or Sleep modes as needed.
@@ III-D2 Dataflow
[…]
@@ III-E Mapping Liquid-S4 to EpochCore
As explained in Section II-A, the Liquid-S4 network shares computational similarities with the S4 network, and all EpochCore innovations for the S4 network are also relevant to the Liquid-S4. One key difference is in the recurrent integration step. Liquid-S4 introduces a time-varying, input-dependent coefficient, which requires the recurrent integration PEs to operate in TRI-MAC mode. The dataflow of the Liquid-S4 network largely mirrors that of the S4 network (as illustrated in Figure 7b). The main difference is that the second-row PEs are preconfigured to execute the TRI-MAC operation. Figure 8b highlights the specific operations of the Liquid-S4 network.
@@ III-F1 CNN/RNN/Transformers
EpochCore can be configured to execute a variety of neural network operations, such as GEMM for CNN, RNN and Transformers, by supporting common dataflows like WS, IS and OS. Figure 9 demonstrates how EpochCore can be utilized for a DNN operation using the OS dataflow. In this configuration, all LIMA-PE are preset to perform the Traditional OS (TOS) MAC operation (Figure 6d) by proper selection of multiplexers (M2, M3, M4 and M6). The inputs and weight vectors are read every cycle and fed to the SA. The input data flows in the West-to-East direction, and the weights flow in the North-to-South direction. At the end of the GEMM operation, the entire output matrix is read out. The IS and WS dataflows can be similarly implemented. Transformers predominantly use GEMM operations as well, and their applications are complementary to those of SSMs. So, transformers can be executed on EpochCore.
@@ III-G EpochCore for other SSMs
Patro et. al [25] provide a taxonomy of recently proposed SSMs, categorizing them based on their structural, gated and recurrent characteristics. Within the structured category, models such as H3 [6] implement two layers of structured SSMs: one performing a shift operation, and another utilizing a diagonal state matrix. The shift-SSM layer in such models can be executed within EpochCore’s SA as a single row of LIMA-PE by including additional sleep PE at the beginning of the row and shifting the row to the right. SSMs in the gated category improve contextualization, as seen in GSS [23]. Mapping GSS models onto EpochCore involves evaluating the associated linear layers by appropriately setting the mode of LIMA-PE, followed by the evaluation of an on-chip non-linear activation function to implement gating. The Mamba model [7, 21], shares key characteristics with Liquid-S4 [13, 12], including the use of input-dependent time-varying state-matrix. Mamba also builds upon the framework introduced in H3 [6] and GSS [23]. In Mamba, the discretized coefficients $\mathbf{\overline{A}}$ and $\mathbf{\overline{B}}$ vary with input, while the discretization step $\triangle t$ effectively serves as a gating mechanism. Supporting Mamba models on EpochCore requires pre-loading the inputs to compute input-dependent discretized coefficients on the host. These updated coefficients must be loaded onto EpochCore as weight matrices, introducing additional overhead for input processing and data transfer. This overhead is roughly $N\times$ the performance cost of S4, where $N$ is the state-map size, making EpochCore highly inefficient. To efficiently support Mamba, two modifications to EpochCore are required. (1) Support variable discretization ($\Delta t$) via additional LIMA-PE rows, (2) new MAC operations in LIMA-PE to support gated element-wise multiplication. These modifications to EpochCore, to make it generic, are part of our future work.
@@ III-H Other Components
A complete neural network model involves additional mathematical operations at the output of each layer. EpochCore integrates on-chip specialized circuitry to support a range of non-linear activation functions (such as SiLU, ReLU, Sigmoid and TanH) and layer normalization, as shown in the EpochCore micro-architecture in Figure 4b. These operations are implemented as custom digital units.
@@ IV Evaluation
In this section, we compare EpochCore accelerator against other SoTA accelerators. The evaluation results apply to both inference and training. Training of neural models requires a forward pass and a backward pass, while inference involves a forward pass. The backward pass includes two GEMM operations for weight updates. In this paper, we evaluate EpochCore for inference i.e. the forward pass using 32-bit fixed-point precision. Improving forward pass performance benefits training as well.
@@ IV-A Methodology
We designed LIMA-PE at the RTL level and synthesized it using Free-PDK45 [31], along with NanGate’s standard cell library. For synthesis, we utilized Cadence’s Genus™ tool [1] to meet a target clock frequency of 700MHz. We created test benches and simulated 1 million cycles using Cadence’s NC-Sim™, generating a Value Change Dump (VCD) file. This file was then fed back into Genus™ to evaluate the quality of results (QoR) using power, area, and performance metrics. To evaluate the LIMA-PE, we compared it against other PEs from the literature. We accounted for differences in technology nodes by applying appropriate technology scaling techniques [28].
@@
To evaluate EpochCore, we developed a custom cycle simulator based on ScaleSim [26] to estimate key metrics such as cycle count, throughput, and energy consumption for various configurations. For evaluating S4 and Liquid-S4 based models, we used model parameters cited by Gu et al  [9]. Our analysis is based on LRA datasets (Table I), where S4-based neural models remain SoTA. Additionally, we created a memory bandwidth simulator to measure continuous bandwidth requirements of EpochCore and other accelerators under LRA workloads. The memory access time, using the same technology i.e. 45 nm as the PE design, was determined using a CACTI-based memory compiler [20].
TABLE III: Comparison of PEs. Traditional-PE, STPU-PE and four types of LIMA-PE for FixedPoint32 and Int8 operation
@@ IV-B LIMA-PE vs other PEs
Sparse-TPU (STPU) [14] was designed to handle GEMM with large sparsity, featuring a PE that supports special modes. When comparing LIMA-PE to STPU-PE, LIMA-PE demonstrates a smaller area footprint for both the 8-bit and 32-bit versions. LIMA-PE is more energy-efficient than STPU-PE in the 8-bit version but consumes more power in the 32-bit version.
@@
The 20-40% reduction in power at the cost of a 30% increase in area is presented in Table III, comparing rows 1 (Trad-PE without clock-gating) and 3 (LIMA-PE with Trad-MAC and with clock-gating). Power consumption for various LIMA-PE operating modes was measured and summarized in Table IV. The Sleep and Pass-Through modes provide substantial power savings, especially when higher bit-precision is used.
@@ IV-C EpochCore vs Systolic Arrays
S4 and Liquid-S4 models involve specialized operations, such as scalar-vector product and recurrent integrations that have sparsity and so they are not well suited to be run on regular SA. To evaluate EpochCore for these models, we compare its performance against other SA-based architectures such as the Sparse-TPU [14] and Boriakoff’s [3] FFT-based SA accelerator. The datasets used in this evaluation were selected based on software evaluations of S4 and Liquid-S4 done by Gu et al [9], which spans vision, audio, and other challenging LRA tasks commonly used to benchmark state-of-the-art model accuracy [21]. For our experiments, we focused on a single Liquid-S4 layer with a state-map size of $N=64$, and varied sequence lengths, $T$ as shown in Table I.
Figure 10: Energy/Inference and Latency/Inference for various LRA datasets shown in Table I for Liquid-S4 model when using Boriakoff-based 1-D SA [3], Sparse 2D SA, and EpochCore.